OpenStudy (erniellyd28):

if a plane can travel 500 miles per hour with the wind and 420 per hour against the wind . find the speed of the plane without a wind and the speed of the wind.

OpenStudy (welshfella):

If x is speed of plane without the wind and y = speed of the wind then we have the system of equations x + y = 500 x - y = 420 solve these 2 equations

OpenStudy (erniellyd28):

250+250=500? 210+210=420